Question 16. You know that put call parity must hold and you observe the following information in the market:

Spot: 195kr

Strike: 180kr

Call premium: 24kr

Put premium: 7kr

Time to maturity: 9 months exactly

(the Call and the Put options have the same underlying security, strike price and maturity date)

  1. What is the risk-free rate?
Answer #1

Put-Call-Parity C + X/(1+rf) = So+P 24kr + 180kr/(1+rf)0.75 = 195kr - 7kr 180kr/(1+rf)0.75= 164kr 180kr/164kr = (1+rf)0.75 1